OpenStudy (anonymous):

If we divide the polynomial x4 + 4x3 + 2x2 + x + 4 by x2 + 3x, what will be the remainder

OpenStudy (usukidoll):

\[x^4+4x^3+2x^2+x+4\] by \[x^2+3x\] have you made any attempts to this problem?

OpenStudy (anonymous):

x4+4x3+2x2+x+4byx2+3x =x4+4x3+2x2+x+4bx2y+3x Combine Like Terms: =x4+4x3+2x2+x+4bx2y+3x =(4bx2y)+(x4)+(4x3)+(2x2)+(x+3x) =4bx2y+x4+4x3+2x2+4x Answer: =4bx2y+x4+4x3+2x2+4x right?

OpenStudy (usukidoll):

whoa man.. direct answers are against openstudy's code of conduct

OpenStudy (usukidoll):

and this is long division... I'm not doing that until the official poster responds

OpenStudy (anonymous):

i have made attempts and none of the answer are there

OpenStudy (usukidoll):

so you had tried the long division method to this problem?

OpenStudy (anonymous):

thats not a direct answer i said it step by step and just to let you know i did this on the other post and they said that they under stood it completely now

OpenStudy (anonymous):

yes sir i have. would you like to see the answer choices

OpenStudy (usukidoll):

yeah I like to see the choices

OpenStudy (anonymous):

A) x2 + 3x B) 4x + 4 C) x2 + x – 2 D) 4 E) 0 That is all of the answer choices

OpenStudy (usukidoll):

ok... so this problem is asking to divide... so we need long division.